Preble Company manufactures one product. Its variable manufacturing overhead is applied to production based on direct labor hours and its standard cost card per unit is as follows:



Direct materials: 5 pounds at $8.00 per pound$40.00
Direct labor: 2 hours at $14 per hour
28.00
Variable overhead: 2 hours at $5 per hour
10.00
Total standard cost per unit$78.00

  

The planning budget for March was based on producing and selling 25,000 units. However, during March the company actually produced and sold 30,000 units and incurred the following costs:

  1. Purchased 160,000 pounds of raw materials at a cost of $7.50 per pound. All of this material was used in production.

  2. Direct laborers worked 55,000 hours at a rate of $15.00 per hour.

  3. The total variable manufacturing overhead for the month was $280,500.


Please answer the following: (show as much work as possible)

1. What raw materials cost would be included in the company’s planning budget for March?

2. What raw materials cost would be included in the company’s flexible budget for March?

3. What is the materials price variance for March?

4.  What is the materials quantity variance for March?

5.  If Preble had purchased 170,000 pounds of materials at $7.50 per pound and used 160,000 pounds in production, what would be the materials price variance for March?

6. if Preble had purchased 170,000 pounds of materials at $7.50 per pound and used 160,000 pounds in production, what would be the materials quantity variance for March?

7. What direct labor cost would be included in the company’s planning budget for March?

8. What direct labor cost would be included in the company’s flexible budget for March?

9. What is the labor rate variance for March?

10. What is the labor efficiency variance for March?

11. What is the labor spending variance for March?

12. What variable manufacturing overhead cost would be included in the company’s planning budget for March?

13. What variable manufacturing overhead cost would be included in the company’s flexible budget for March?

14. What is the variable overhead rate variance for March?

15. What is the variable overhead efficiency variance for March?
Answer #2

Planning budgetFLexible budgetActual
Level of activities250003000030000
Direct material1000000
12000001200000

Direct Labor700000840000825000F15000
V MOH250000300000280500
F19500

Variance analysis

Price/ rate variance
Quantity/ Efficiency Variance
Direct material(7.5-8)*160000
80000F(160000-150000)*880000U
Direct Labor
55000
55000U(55000-60000)*1470000F
MOH VMOH
55000U F
25000F

Total: Labor: 15000F

MOH: 19500 F


answered by: Nguyet
Answer #1

1. Raw materials cost in planning budget: 40*25000=10,00,000

2. Raw materials cost in flexible budget: 7.50*160000=12,00,000

3.Material cost variance: 10,00,000-12,00,000=-2,00,000

4. Material quantity variance: (159990-125000)*8= 279920

5. Material cost variance: 10,00,000-1275000=-275000

6. Material quantity variance: (159990-125000)*8= 279920

7. Direct Labor: 28*25000=700000

8. Direct Labor: 55000*15= 825000

9. Labor rate variance:55000*15-55000*14=55000

10. Labor efficiency variance= (2- 1.833)*14=2.333

11.Labor spending variance= 76.85-78=-1.15

12 variable manufacturing overhead : 10*25000=250000

13 variable manufacturing overhead : 280500

15variable overhead efficiency variance: (55000-50000)*5=25000